OpenStudy (anonymous):

Select the equation of a line that is parallel to the line on the graph and passes through the point (2, 3). (there is a graph) y = 2x + 4 y = - 2x + 2 y = - 1 over 2 x + 2 y = - 1 over 2 x + 4

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

It'd be easier to plug in the vales and see which statement is true.

OpenStudy (anonymous):

okay I'm trying to get the graph up so you can see it one second please

OpenStudy (anonymous):

there is the graph

OpenStudy (luigi0210):

Oh, nevermind my first statement then. What we have to do is find the slope of the graph. Do you know it?

OpenStudy (anonymous):

i've looked and I think c or d.

OpenStudy (luigi0210):

Yup, so the only way I know how to do it is using the point-slope formula: \[\LARGE y-y_{1}=m(x-x_{1})\] Do know this formula?

OpenStudy (anonymous):

yes I have seen that

OpenStudy (luigi0210):

So m is the slope, and x and y are the points: So just simplify this and you should get the right answer: \[\LARGE y-3=-\frac{1}{2}(x-2)\]

OpenStudy (anonymous):

how do I simplify that? I always forget how to do that.

OpenStudy (luigi0210):

Just distribute the 1/2 first.

OpenStudy (anonymous):

so it'd be something like 3 1/2? Im really sorry Im not to good at stuff like this.

OpenStudy (luigi0210):

Not exactly, think of it this way: \[\LARGE y-3=(-\frac{1}{2}*x)-(\frac{1}{2}*(-2))\]

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

i think D is right. do u think so?

OpenStudy (luigi0210):

Yup, it would be D.

OpenStudy (anonymous):

cool thanks again.

OpenStudy (luigi0210):

Just as a heads up for future reference. You can give medals by hitting the "Best Response" button. Someone are picky about that, just letting you know now so people don't bug you about it later.

OpenStudy (anonymous):

okay thanks

OpenStudy (luigi0210):

Yup!